# Physics-informed DeepCT: Sinogram Wavelet Decomposition Meets Masked Diffusion

### v2

@@ IV-C Reconstruction Experiments
IV-D Profile Lines Analysis
In this study, we evaluated the edge-preserving performance of different reconstruction methods by comparing the generated contours with the ground truth contours. As shown in Fig. 8, the contours generated by FBP, FBPConvNet, HDNet, and SWORD exhibit poor alignment with the ground truth, with noticeable deviations in edge details. In contrast, the contours generated by GMSD and SWARM exhibit a closer match to the ground truth, further highlighting the advantage of diffusion models in preserving edge details. Among these methods, the contours generated by SWARM are the most accurate, particularly in regions with significant gray-scale variations, where the consistency between SWARM generated contours and the actual data is most evident. These results further validate the superiority and robustness of the SWARM method in edge preservation.
@@ IV-D Profile Lines Analysis
IV-E Ablation and Generalization Study
1) Mask Ratio Analysis: To systematically evaluate the impact of different mask sparsity ratios on image reconstruction performance, ablation experiments were conducted with mask ratios of 0%, 10%, 20%, 30%, 50%, and 70%, as illustrated in Fig. 9. As the masking ratio increases, both PSNR and SSIM initially improve, plateau, and then decline, while MSE demonstrates a typical “U-shaped” curve-decreasing initially and rising again beyond a certain point. This trend suggests that extremely low or high masking ratios hinder the model’s ability to effectively infer missing information, resulting in suboptimal reconstruction quality. In contrast, a moderate masking level (e.g., 20%) introduces a beneficial degree of sparsity that stimulates the model’s prior learning capacity, achieving an optimal balance between reconstruction quality and robustness. These findings further validate the hypothesis proposed in the methodological motivation: introducing an appropriate level of random masking enhances the model’s ability to complete unobserved regions, thereby improving overall reconstruction performance. This provides clear experimental evidence and guidance for selecting optimal mask ratios in future applications.
2) Different Components in SWARM: SWARM employs a dual-diffusion model for iterative reconstruction. During this process, SRM is trained using a random mask strategy, while SHD is trained based on random high-frequency components from the sinogram. The quantitative analysis in Table IV demonstrates that SWARM significantly improves reconstruction quality. Fig. 10 further illustrates the impact of the model combination. Qualitative results show that the cascade of SRM and SHD in iterative reconstruction effectively leverages their complementary advantages, leading to a significant enhancement in image reconstruction quality. This outcome validates the effectiveness of our approach, showing that combining the strengths of both models achieves superior reconstruction performance.
3) Effectiveness of Random Masking and High-Frequency of Sinogram: As shown in Table V and Fig. 11, we studied the effects of the random masking strategy and randomness based on high-frequency components in the wavelet domain from both quantitative metrics and visual perspectives. NMS represents the non-masked condition for the sinogram, while NRH represents the non-random wavelet condition for the sinogram. Generalization experiments conducted on the CIRS phantom data further confirmed that the incorporation of random masking not only significantly improved the model’s generalization ability but also enhanced its robustness, thereby validating the effectiveness of the proposed method. Additionally, the strategy of randomly selecting high-frequency components as network training inputs demonstrated good stability, effectively boosting the model’s generalization capability and improving overall reconstruction efficiency.
TABLE IV: Ablation Study of Two-Stage Model on AAPM Challenge Data.
@@ IV-E Ablation and Generalization Study
TABLE V: Generalization Study of Random Masking & Wavelet HF.
@@
IV-F Computational Complexity
As shown in Table VI, the comparison of the computational performance of various methods under 60 views, FBPConvNet demonstrates the optimal computational efficiency. However, FBPConvNet relies on separate training and solving for each scanning view. This angular dependency significantly increases the complexity of the sampling process and the training cost. In particular, it limits the practicality of the method in multi-view or dynamic imaging tasks. In contrast, although methods such as GMSD, SWORD, and SWARM have relatively higher time complexity, they only require training a single model for multi-angle reconstruction tasks. This significantly enhances the flexibility and generalization ability of these methods in application scenarios, especially in imaging environments with frequent angular changes or limited sampling conditions.
TABLE VI: Quantitative Evaluation of Reconstruction Time.
Methods FBPConvNet HDNet GMSD SWORD SWARM
Time 0.15s 1.27s 461.58s 1214.21s 1290.43s
V Discussion and Conclusion

### v3

@@ 4.3.3 Dental Arch Reconstruction Results
4.4 Profile Lines Analysis
In this study, we further evaluated the reconstruction accuracy of different methods through profile analysis. This analysis quantitatively assesses the reconstruction quality by comparing the signal intensity distributions along specific cross-sectional regions between the reconstructed images and the ground truth. As shown in Fig. 7, the signal intensity profiles of FBP, FBPConvNet, HDNet, and SWORD exhibit noticeable deviations from the ground truth, particularly in regions with sharp intensity transitions, indicating their limitations in accurately restoring structural boundaries. In contrast, GMSD and SWARM demonstrate much higher consistency with the ground truth, reflecting their superior capability in preserving local structures and intensity variations. Notably, SWARM achieves the closest numerical correspondence to the ground truth in both smooth and high-contrast regions, suggesting that it not only maintains structural integrity but also delivers superior quantitative accuracy—an attribute that is particularly crucial for medical diagnosis. These results further validate the robustness and reliability of the proposed SWARM method in reconstructing clinically relevant image details.
@@ 4.4 Profile Lines Analysis
4.5 Ablation and Generalization Study
4.5.1 Mask Ratio Analysis
To systematically evaluate the impact of different mask sparsity ratios on image reconstruction performance, ablation experiments were conducted with mask ratios of 0%, 10%, 20%, 30%, 50%, and 70%, as illustrated in Fig. 8. As the masking ratio increases, both PSNR and SSIM initially improve, plateau, and then decline, while MSE demonstrates a typical “U-shaped” curve-decreasing initially and rising again beyond a certain point. This trend suggests that extremely low or high masking ratios hinder the model’s ability to effectively infer missing information, resulting in suboptimal reconstruction quality. In contrast, a moderate masking level (e.g., 20%) introduces a beneficial degree of sparsity that stimulates the model’s prior learning capacity, achieving an optimal balance between reconstruction quality and robustness. These findings further validate the hypothesis proposed in the methodological motivation: introducing an appropriate level of random masking enhances the model’s ability to complete unobserved regions, thereby improving overall reconstruction performance. This provides clear experimental evidence and guidance for selecting optimal mask ratios in future applications.
4.5.2 Different Components in SWARM
SWARM employs a dual-diffusion model for iterative reconstruction. During this process, SRM is trained using a random mask strategy, while SHD is trained based on random high-frequency components from the sinogram. The quantitative analysis in Table 4 demonstrates that SWARM significantly improves reconstruction quality. Fig. 9 further illustrates the impact of the model combination. Qualitative results show that the cascade of SRM and SHD in iterative reconstruction effectively leverages their complementary advantages, leading to a significant enhancement in image reconstruction quality. This outcome validates the effectiveness of our approach, showing that combining the strengths of both models achieves superior reconstruction performance.
4.5.3 Effectiveness of Random Masking and High-Frequency of Sinogram
As shown in Table 5 and Fig. 10, we studied the effects of the random masking strategy and randomness based on high-frequency components in the wavelet domain from both quantitative metrics and visual perspectives. NMS represents the non-masked condition for the sinogram, while NRH represents the non-random wavelet condition for the sinogram. Generalization experiments conducted on the CIRS phantom data further confirmed that the incorporation of random masking not only significantly improved the model’s generalization ability but also enhanced its robustness, thereby validating the effectiveness of the proposed method. Additionally, the strategy of randomly selecting high-frequency components as network training inputs demonstrated good stability, effectively boosting the model’s generalization capability and improving overall reconstruction efficiency.
@@ 4.5.3 Effectiveness of Random Masking and High-Frequency of Sinogram
Table 4: Ablation Study of Two-Stage Model on AAPM Challenge Data.
@@
Table 5: Generalization Study of Random Masking & Wavelet HF.
@@
5 Discussion and Conclusion